# Matching Non-Identical Objects

### v1

@@ 3.2 Technical challenges
Generalization from image matching to non-identical object matching introduces ambiguities and perturbations into matching. First, because there is no “correct matching” between two non-identical objects (e.g., a dog and a cat), one cannot resort to a simple supervised learning framework. There is not clear measure of a good matching that can be used for computing a loss to minimize. Second, domain shift by common corruptions or image style significantly changes the low-level features. To our knowledge, existing matching methods only use low-level features. such as colors, edges, and textures. As a consequence, these methods are susceptible to the domain shift, as shown in Fig. 2. The first challenge suggests that it would be better to somehow robustifies pretrained image matching models instead of training a model from scratch. The second challenge indicates the necessity of introducing semantic information, which is known to be more robust against domain shifts.
@@ 4 Proposed Method
In this section, we present a simple plug-and-play weighting module that can turn various sparse matching models to work for non-identical object matching. Most sparse matching models, such as [30, 17, 3, 23], consist of a pair of a keypoint detector and a feature matcher, and our module process the keypoints before feeding them to the feature matcher. Particularly, given a pair of images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$, keypoint detectors output collections of keypoints $D_{A}=\{(\bm{p}_{i}^{A},\bm{d}_{i}^{A})\}_{i},D_{B}=\{(\bm{p}_{i}^{B},\bm{d}_{i}^{B})\}_{i}$. The proposed module then computes weighted descriptors $\widetilde{\bm{d}}_{i}^{A},\widetilde{\bm{d}}_{i}^{B}$ for each $\bm{d}_{i}^{A},\bm{d}_{i}^{B}$, respectively. Lastly, the feature matcher performs matching using $\widetilde{D}_{A}=\{(\bm{p}_{i}^{A},\widetilde{\bm{d}}_{i}^{A})\}_{i},\widetilde{D}_{B}=\{(\bm{p}_{i}^{B},\widetilde{\bm{d}}_{i}^{B})\}_{i}$.
@@ 4.1 Weighting module
The proposed weighting module uses an object detector and visual explanation model. In this study, we used YOLOv7 [41] and Grad-CAM [32]. Let $\bm{x}\in{\mathcal{X}}$ be an image, and let $D=\{(\bm{p}_{i},\bm{d}_{i})\}_{i=1}^{n}$ be the collection of $n$ keypoints of $\bm{x}$.
First, the object detector is applied to $\bm{x}$ and detects the objects with their classes. For each of the detected objects (say, the $k$-th object), the visual explanation model gives pixel-wise class-activation scores, or a heatmap $H_{k}$. Let $H_{k,\bm{p}_{i}}\in[0,1]$ be the heatmap score of $H_{k}$ at $\bm{p}_{i}$. Then, the heatmap scores at $\bm{p}_{i}$ are aggregated as $H_{\bm{p}_{i}}:=\max_{k}H_{k,{\bm{p}_{i}}}$. Finally, the weighted descriptor $\tilde{d}_{i}$ is computed by
@@
We empirically found that the addition of base value 1 and the normalization by the maximum value is important. For example, the simplest approach $\widetilde{\bm{d}}_{i}=H_{\bm{p}_{i}}\bm{d}_{i}\in[0,1]^{d}$ ends up with a scarce of matching because of the heatmap scores on the keypoints on background tend to be too low, i.e., $H_{\bm{p}_{i}}\approx 0$. Although such keypoints will not likely appear in the matching results, they seem to help the keypoints on objects to be matched. If one instead uses $\widetilde{\bm{d}}_{i}=(1+H_{\bm{p}_{i}})\bm{d}_{i}\in[1,2]^{d}$, this also affects negatively. We consider that this is because of the softmax operations used in the subsequent feature matching. We discuss the consequence of the weighting of descriptors next.
@@ 4.2 Effect of weights in feature matcher
We now illustrate how our weighting module affects the subsequent feature matching. We take LightGlue [17] as an example, but a similar discussion holds for other feature matchers. For simplicity, let $\alpha_{i}$ be the weight in Eq. 5, i.e., $\widetilde{\bm{d}}_{i}=\alpha_{i}\bm{d}_{i}$. The feature matcher takes as input the keypoints $\widetilde{D}_{A}=\{(\bm{p}_{i}^{A},\widetilde{\bm{d}}_{i}^{A})\}_{i},\widetilde{D}_{B}=\{(\bm{p}_{i}^{B},\widetilde{\bm{d}}_{i}^{B})\}_{i}$. The vectors $\widetilde{\bm{d}}_{i}^{A}=\alpha_{i}^{A}\bm{d}_{i}^{A},\widetilde{\bm{d}}_{i}^{B}=\alpha_{i}^{B}\bm{d}_{i}^{B}$ are the $i$-th weighted descriptors of image $\bm{x}_{A},\bm{x}_{B}$, respectively. Below, symbols with and without $\widetilde{\,\cdot\,}$ present those with and without weighting.
@@
where $I\in\{A,B\}$. Then, the self-attention score $a_{ij}^{I}$ between $i$-th and $j$-th keypoints of image $\bm{x}_{I}$ is calculated as follows. [^1] To be precise, this score should be considered as the pre-activation score because a softmax operation will be subsequently applied. The terminology of this paper follows the related studies, e.g., [17].
@@
$=(\alpha_{i}^{I}\alpha_{j}^{I})\cdot a_{ij}^{I}.$ (10)
where $\mathcal{R}:{\mathbb{R}}^{2}\to{\mathbb{R}}^{d\times d}$ denotes the rotary-encoding [33]. Therefore, our weighting by heatmap scores leads to weighting of attention scores. The self-attention scores reflect the internal relationship of objects and background, thereby extracting better feature representations in the subsequent steps of the feature matcher. Roughly speaking, nothing that $\alpha_{i}^{I}\in[0.5,1.0]$ for all $i$ and $I\in\{A,B\}$, we have $\alpha_{i}^{I}\alpha_{j}^{I}\approx 1$ between the keypoints on objects, $\alpha_{i}^{I}\alpha_{j}^{I}\approx 0.5$ between one on object and the other on background, and $\alpha_{i}^{I}\alpha_{j}^{I}\approx 0$ for between those on the background. This is reasonable because we want nice feature representations for object keypoints for nice matching. Nevertheless, such nice representations should be obtained by taking into account the background conditions, e.g., to compensate for the overall lighting conditions. If one uses $\alpha_{i}^{I}=H_{\bm{p}_{i}}$ as discussed earlier, i.e., $H_{\bm{p}_{i}}\approx 0$ for background keypoints, the feature representations cannot take into account the background.
Similarly, our weighting affects the cross attention scores. The cross-attention score $a_{ij}^{AB}$ between the $i$-th keypoint of image $\bm{x}_{A}$ and the $j$-th keypoint of image $\bm{x}_{B}$ is computed from the key vectors.
@@
As discussed earlier, the value of $\alpha_{i}^{A}\alpha_{j}^{B}$ are generally large between the keypoints on objects. Clearly, this encourages the matching between objects.
@@ 5 Experiments
We now demonstrate the effectiveness of our method in non-identical object matching tasks. Particularly, we first present qualitative results on various pairs of non-identical (but similar) objects. We then quantitatively evaluate the proposed methods on classical image matching tasks but under common corruptions.
@@
In the sparse matching framework, the image matching task is performed by a keypoint detector and a feature matcher. For the former, we adopted pre-trained SuperPoint [7], extracting 2,048 keypoints. For the latter, the extracted keypoints are matched using either of pretrained SGMNet [3], GlueStick [23], and the state-of-the-art method, LightGlue [17], with and without the proposed weighting module. We additionally tested a detector-free dense matcher, LoFTR [36]. We downloaded the pretrained models from their official repositories.222 SGMNet: https://github.com/vdvchen/SGMNet
GlueStick: https://github.com/cvg/GlueStick
LightGlue: https://github.com/cvg/LightGlue
LoFTR: https://github.com/zju3dv/LoFTR
All the experiments are conducted on a computer with an NVIDIA GeForce RTX 3090 GPU and 24GB memory.
@@ 5.1 Matching Non-Identical Objects
We examined four cases: (i) same class (i.e., $y_{A}=y_{B},D_{A}=D_{B}$), (ii) different class (i.e., $y_{A}\neq y_{B},D_{A}=D_{B}$), (iii) another domain (i.e., $o_{A}\equiv o_{B}$, $D_{A}\neq D_{B}$), and (iv) same appearance (i.e., $o_{A}=o_{B}$). For all the cases, we have perspective difference (i.e., $P_{A}\neq P_{B}$). Due to the page number restriction, we only present a limited number of matching examples here, but more results can be found in the supplementary material. Figure 4 shows the results of the matching results with and without the proposed method in the four cases. As can be seen in Fig. 4(a), a combination of SuperPoint and LightGlue cannot distinguish between objects and backgrounds or correspond to the same structure of animals. In contrast, Fig. 4(b) shows that the introduction of the proposed method significantly improves the results, reducing unreasonable matchings (e.g., human and racket, animal eye and mouth) and increasing reasonable part-to-part matchings. Particularly, for the cases (i) increased the number of matches between rackets and decreased the mismatches between people, (ii) corresponded to the same parts of animals, (iii) matched rear wheels with domain gaps, and (iv) significantly increased the number of matches despite changes in background and object angles. Notably, structural features, such as eyes and noses, are successfully matched between different classes. We consider that the heatmaps of these animals gave similar importance to the eyes or mouth, thus increasing the number of matches between the same parts of the body. In addition, the results for the same class and appearance showed a decrease in mismatches with the background, which means that the background and foreground had relatively lower similarity.
@@
Figure 5 presents the results with different feature matchers, showing that the proposed method boosts all of them to match corresponding parts of the cheetah and Siberian husky.
Figure 5: The results with different feature matchers [3, 23, 17]. SuperPoint [7] is used for the keypoint detection. The introduction of the proposed module increases the matching of eyes and noses between the cheetah and Siberian husky and decreases mismatching in the background.

### v3

@@ 3.2 Technical challenges
Generalization from image matching to non-identical object matching introduces ambiguities and perturbations into matching. First, because there is no “correct matching” between two non-identical objects (e.g., a dog and a cat), one cannot resort to a simple supervised learning framework. There is no clear measure of a good matching that can be used for computing a loss to minimize. Second, domain shifts by common corruptions or image styles significantly change the low-level features. To our knowledge, existing matching methods only use low-level features such as colors, edges, and textures. As a consequence, these methods are susceptible to the domain shift, as shown in Fig. 2. The first challenge suggests that it would be better to somehow robustify pretrained image matching models instead of training a model from scratch. The second challenge indicates the necessity of introducing semantic information, which is known to be more robust against domain shifts.
@@ 4 Proposed Method
In this section, we present a simple plug-and-play weighting module that can turn various sparse matchers to work for non-identical object matching. Most sparse matchers, such as [49, 27, 4, 38], consist of a pair of a keypoint detector and a feature matcher, and our module process the keypoints before feeding them to the feature matcher. Particularly, given a pair of images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$, The keypoint detector outputs collections of keypoints $D_{A}=\{(\bm{p}_{i}^{A},\bm{d}_{i}^{A})\}_{i},D_{B}=\{(\bm{p}_{i}^{B},\bm{d}_{i}^{B})\}_{i}$. The proposed module then computes weighted descriptors $\widetilde{\bm{d}}_{i}^{A},\widetilde{\bm{d}}_{i}^{B}$ for each $\bm{d}_{i}^{A},\bm{d}_{i}^{B}$, respectively. Lastly, the feature matcher performs matching using $\widetilde{D}_{A}=\{(\bm{p}_{i}^{A},\widetilde{\bm{d}}_{i}^{A})\}_{i},\widetilde{D}_{B}=\{(\bm{p}_{i}^{B},\widetilde{\bm{d}}_{i}^{B})\}_{i}$.
@@ 4.1 Weighting module
The proposed weighting module uses an object detector and visual explanation model. This study used YOLOv7 [62] and Grad-CAM [51]. Let $\bm{x}\in{\mathcal{X}}$ be an image, and let $D=\{(\bm{p}_{i},\bm{d}_{i})\}_{i=1}^{n}$ be the collection of $n$ keypoints of $\bm{x}$.
First, the object detector is applied to $\bm{x}$ and detects the objects with their classes. For each of the detected objects (say, the $k$-th object), the visual explanation model gives pixel-wise class-activation scores or a heatmap $H_{k}$. Let $H_{k,\bm{p}_{i}}\in[0,1]$ be the heatmap score of $H_{k}$ at $\bm{p}_{i}$. Then, the heatmap scores at $\bm{p}_{i}$ are aggregated as $H_{\bm{p}_{i}}:=\max_{k}H_{k,{\bm{p}_{i}}}$. Finally, the weighted descriptor $\tilde{\bm{d}}_{i}$ is computed by
@@
We empirically found that the addition of base value 1 and the normalization by the maximum value is important. For example, the simplest approach $\widetilde{\bm{d}}_{i}=H_{\bm{p}_{i}}\bm{d}_{i}\in[0,1]^{d}$ ends up with a scarcity of matching because of the heatmap scores on the keypoints on background tend to be too low, i.e., $H_{\bm{p}_{i}}\approx 0$. Although such keypoints will not likely appear in the matching results, they seem to help the keypoints on objects to be matched. If one instead uses $\widetilde{\bm{d}}_{i}=(1+H_{\bm{p}_{i}})\bm{d}_{i}\in[1,2]^{d}$, this also affects negatively. We consider that this is because of the softmax operations used in the subsequent feature matching. We discuss the consequence of the weighting of descriptors next.
@@ 4.2 Effect of weights in feature matcher
We now illustrate how our weighting module affects the subsequent feature matching. We take LightGlue [27] as an example, but a similar discussion holds for other feature matchers. For simplicity, let $\alpha_{i}$ be the weight in Eq. 5, i.e., $\widetilde{\bm{d}}_{i}=\alpha_{i}\bm{d}_{i}$. The matcher takes as input the keypoints $\widetilde{D}_{A}=\{(\bm{p}_{i}^{A},\widetilde{\bm{d}}_{i}^{A})\}_{i},\widetilde{D}_{B}=\{(\bm{p}_{i}^{B},\widetilde{\bm{d}}_{i}^{B})\}_{i}$. The vectors $\widetilde{\bm{d}}_{i}^{A}=\alpha_{i}^{A}\bm{d}_{i}^{A},\widetilde{\bm{d}}_{i}^{B}=\alpha_{i}^{B}\bm{d}_{i}^{B}$ are the $i$-th weighted descriptors of image $\bm{x}_{A},\bm{x}_{B}$, respectively. Below, symbols with and without $\widetilde{\,\cdot\,}$ present those with and without weighting.
@@
where $I\in\{A,B\}$. Then, the self-attention score $a_{ij}^{I}$ between $i$-th and $j$-th keypoints of image $\bm{x}_{I}$ is calculated as [^1] To be precise, this score should be considered as the pre-activation score, and the softmax operation will be subsequently applied. The terminology of this paper follows the related studies, e.g., [27].
@@
$=(\alpha_{i}^{I}\alpha_{j}^{I})\cdot a_{ij}^{I},$ (10)
where $\mathcal{R}:{\mathbb{R}}^{2}\to{\mathbb{R}}^{d\times d}$ denotes the rotary-encoding [53]. Therefore, our weighting by heatmap scores leads to the weighting of attention scores. The self-attention scores reflect the internal relationship of objects and background, thereby extracting better feature representations in the subsequent steps of the feature matcher. Roughly speaking, nothing that $\alpha_{i}^{I}\in[0.5,1.0]$ for all $i$ and $I\in\{A,B\}$, we have $\alpha_{i}^{I}\alpha_{j}^{I}\approx 1$ between the keypoints on objects, $\alpha_{i}^{I}\alpha_{j}^{I}\approx 0.5$ between one on object and the other on background, and $\alpha_{i}^{I}\alpha_{j}^{I}\approx 0$ for between those on the background. This is reasonable because we want nice feature representations for object keypoints for nice matching. Nevertheless, such nice representations should be obtained by taking into account the background conditions, e.g., to compensate for the overall lighting conditions. If one uses $\alpha_{i}^{I}=H_{\bm{p}_{i}}$ as discussed earlier, i.e., $H_{\bm{p}_{i}}\approx 0$ for background keypoints, the feature representations cannot take into account the background.
Similarly, our weighting affects the cross-attention score $a_{ij}^{AB}$ between the $i$-th keypoint of image $\bm{x}_{A}$ and the $j$-th keypoint of image $\bm{x}_{B}$, computed from the key vectors.
@@
As discussed earlier, the value of $\alpha_{i}^{A}\alpha_{j}^{B}$ is generally large between the keypoints on objects. Clearly, this encourages the matching between objects.
4.3 Class Similarity Adapter
Figure 4: Class Similarity Adapter converts class labels from an object detector into embeddings using the CLIP [42] text encoder and compares them with cosine similarity to identify semantically similar object pairs. The GradCAM [51] heatmap of object detector YOLOv7 [62] generated for the chosen object pair is used for our weighting method, enhancing matching with multiple objects. Here, we used the keypoint detector SuperPoint [10] and the feature matcher LightGlue [27].
We propose the Class Similarity Adapter to address scenarios where multiple objects are present in an image. Our weighting module uses all heatmaps of objects detected by the object detector for weighting. However, this can needlessly emphasize irrelevant objects, leading to scattered matching from one to multiple objects. As shown in  Fig. 4 (bottom right), relying solely on visual information leads to matching based on similarities in texture, color, and shape, where similar black features such as the nose of the dog and the ear of the cat, are incorrectly matched, causing correspondences to scatter across multiple animals.
To resolve such confusion in the visual space, we again leverage semantic information given by the object detector. Particularly, we map the class labels of detected objects to text embedding space and determine the object pair to be matched based on the cosine similarity between embedding vectors. Suppose that an image $\bm{x}_{A}$ contains $n_{A}$ objects, and let $Y_{A}=\{y_{i}^{A}\}_{i=1}^{n_{A}}$ be their class labels. We define $Y_{B}=\{y_{i}^{B}\}_{i=1}^{n_{B}}$ similarly for the other image $\bm{x}_{B}$. Let $\mathcal{T}$ be a pretrained text encoder that maps a label $y$ to an embedding vector $\mathcal{T}(y)$. Particularly, we used the CLIP text encoder [42]. The Class Similarity Adapter determines the object pair $(i,j)$ to be matched based on the cosine similarity between the embedding vectors as follows.
$(y_{i}^{A},y_{j}^{B})=\underset{y^{A}\in Y_{A},~y^{B}\in Y_{B}}{\text{argmax}}\;\frac{\mathcal{T}(y^{A})^{\top}\mathcal{T}(y^{B})}{\|\mathcal{T}(y^{A})\|\|\mathcal{T}(y^{B})\|},$ (12)
where $\|\,\cdot\,\|$ denotes the $L_{2}$ norm of vector. Then, the weighting module only uses the Grad-CAM heatmaps of the $i$-th object in $\bm{x}_{A}$ and the $j$-th object in $\bm{x}_{B}$. Based on this result, the weighting module uses the corresponding heatmaps for weighting the descriptors. Figure 4 shows this pipeline.
@@ 5 Experiments
We demonstrate the effectiveness of our method in non-identical object matching tasks, presenting qualitative results on various pairs of non-identical (but similar) objects, followed by quantitative evaluation on classical image matching tasks but under common corruptions [20].
@@
In the sparse matching framework, the image matching task is performed by a keypoint detector and a feature matcher. For the former, we adopted pre-trained SuperPoint [10], extracting 2,048 keypoints. For the latter, the extracted keypoints are matched using either pretrained SGMNet [4], GlueStick [38], and the state-of-the-art method, LightGlue [27], with and without the proposed weighting module. We additionally tested a detector-free dense matcher, LoFTR [56] and Efficient LoFTR [64]. We downloaded the pretrained models from their official repositories (details provided in the supplementary material). All experiments are conducted on a computer with an NVIDIA GeForce RTX 3090 GPU and 24GB memory.
@@ 5.1 Matching Non-Identical Objects
We examined four cases: (i) same class (i.e., $y_{A}=y_{B},D_{A}=D_{B}$), (ii) class discrepancy (i.e., $y_{A}\neq y_{B},D_{A}=D_{B}$), (iii) domain shift (i.e., $o_{A}\equiv o_{B}$, $D_{A}\neq D_{B}$), and (iv) same appearance (i.e., $o_{A}=o_{B}$). For all the cases, we have perspective difference (i.e., $P_{A}\neq P_{B}$). Due to the page number restriction, we only present a limited number of matching examples here, but more results can be found in Sec. A in the supplementary material. Figure 5 shows the matching results with and without the proposed method in the four cases. As can be seen in Fig. 5(a), a combination of SuperPoint and LightGlue cannot distinguish between objects and backgrounds or correspond to the same structure of animals. In contrast, Fig. 5(b) shows that the introduction of the proposed method significantly improves the results, reducing unreasonable matchings (e.g., human and racket, animal eye and mouth) and increasing reasonable part-to-part matchings. Particularly, for the cases (i) increased the number of matches between rackets and decreased the mismatches between people, (ii) corresponded to the same parts of animals, (iii) matched front wheels with domain gaps, and (iv) significantly increased the number of matches despite changes in background and object angles. Notably, structural features, such as eyes and noses, are successfully matched between different classes. We consider that the heatmaps of these animals gave similar importance to the eyes or mouth, thus increasing the number of matches between the same parts of the body. In addition, the results for the same class and appearance showed a decrease in mismatches with the background, which means that the background and foreground had relatively lower similarity.
@@
Figure 6 presents the results with other matchers, showing that our method boosts all of them to match corresponding parts between different animals.
Table 1: The pose accuracy (AUC) at the maximum angular error of $20\tcdegree$ of the relative pose estimation from image pairs MegaDepth-1500 [56] under common corruptions [20]. The left side shows both images corrupted with the same type of corruption, while the right side shows only one of the input images corrupted and the other as a clean image. Our method largely improves LightGlue (LG) [27] and GlueStick (GS) [38] for most categories and the average AUC. The dense matchers, LoFTR [56] and Efficient LoFTR (ELoFTR) [64], are shown for reference. Our method improves the robustness of sparse matchers with a slight reduction in clean accuracy and runtime.
[…]
keypoint detector : SuperPoint dense matcher keypoint detector : SuperPoint dense matcher
LG LG+Ours GS GS+Ours LoFTR ELoFTR LG LG+Ours GS GS+Ours LoFTR ELoFTR
None (Clean) 80.61 78.42 78.04 75.08 80.93 83.48 80.61 78.42 78.04 75.08 80.93 83.48
Gaussian Noise 43.09 53.27 45.67 52.24 34.97 35.71 27.54 41.01 41.93 44.98 33.02 36.56
Shot Noise 43.41 53.82 46.10 51.72 40.00 40.36 32.10 42.35 41.75 43.82 36.53 39.52
Impulse Noise 44.98 50.11 44.91 50.05 37.73 39.89 35.95 43.67 40.97 42.91 34.92 38.63
Defocus Blur 32.69 48.35 45.58 49.23 57.22 54.97 18.25 32.13 23.16 30.10 49.40 41.97
Frosted Glass Blur 33.37 47.88 47.40 49.94 54.48 47.09 34.25 44.80 42.96 46.21 54.53 47.52
Motion Blur 42.12 53.63 44.76 51.84 55.60 48.12 50.07 53.10 44.27 52.69 52.81 45.10
Zoom Blur 24.40 31.21 22.87 30.97 23.08 22.54 34.67 40.04 31.18 37.11 31.09 28.49
Snow 31.51 30.33 25.26 23.21 31.85 37.89 56.24 58.15 47.22 48.53 49.15 54.56
Frost 32.24 31.71 26.39 24.22 19.23 25.49 62.20 64.87 54.34 57.26 43.17 45.18
Fog 70.99 73.49 71.19 73.89 67.10 67.74 76.96 78.40 75.18 74.87 74.06 75.36
Brightness 75.48 75.08 73.03 70.84 75.45 79.83 77.14 76.98 76.78 75.06 78.41 81.26
Contrast 39.50 38.47 39.47 42.61 53.22 59.78 43.09 45.17 40.44 43.53 39.66 52.76
Elastic Transform 54.78 66.21 60.37 64.36 58.56 60.50 64.93 68.34 68.02 69.92 68.95 71.49
Pixelate 67.94 68.24 65.01 64.81 74.77 77.27 66.80 68.43 67.16 65.46 76.21 78.45
JPEG Compression 29.54 36.97 39.66 44.70 59.05 59.80 47.41 53.82 60.39 60.12 68.14 71.70
Average 44.40 50.58 46.51 49.64 49.49 50.47 48.51 54.08 50.38 52.84 52.67 53.90
Time [ms/pair] 46.67 68.43 105.72 128.18 312.54 191.17 45.98 68.37 105.34 127.86 305.01 192.53
@@ 5.2 Robust Image Matching
We next quantitatively evaluate the proposed method on a standard image matching task but with common corruptions. To our knowledge, we present the first evaluation of the robustness of image matching under common corruptions. We address two cases: (i) both input images are corrupted, and (ii) only one of them is corrupted. The former shows the robustness against corruption, and the latter shows the robustness against domain shift (i.e., clean image domain v.s. corrupted image domain).
Figure 6: The results with different matchers [4, 38, 27]. SuperPoint [10] is used for the keypoint detection. The introduction of our module increases the matching of eyes and noses between the cheetah and husky and decreases mismatching in the background.
Dataset.
We use image pairs from the MegaDepth-1500 [56] test set as the basis for various experiments. MegaDepth [25] dataset consists of 3D reconstructed data from one million images for two popular phototourism destinations, with camera poses and depth maps computed via COLMAP [50] and refined as ground truth. Common Corruptions [20] simulates the 15 types of corruptions that occur when images are captured. The severity level of corruption is set to 5, which is the most challenging condition.
Metric.
We follow the standard evaluation of image matching based on relative pose estimation (e.g., see [27]). From the matching results, the relative pose between two images is computed with RANSAC [13] as follows. First, an essential matrix representing the positional transformation between the images is calculated, which is then decomposed into a rotation and a translation. The camera pose error is computed by the maximum angular error, and the area under the cumulative error curve (AUC) at $20\tcdegree$ is calculated.
Robustness against common corruptions.
The left side of Tab. 1 shows the AUC of matching when input images are both corrupted. The results of the dense matching method (i.e., LoFTR and Efficient LoFTR) are given for reference. As the results show, the proposed method makes the LightGlue and GlueStick robust against most types of common corruptions (roughly 5% to 10% increase) with a slight increase in the runtime. The average AUC even exceeds that of Efficient LoFTR. The matching methods benefit from the proposed method, particularly when images are corrupted by noises and blurs. We observe AUC decrease in a few types of corruptions, such as Snow and Frost, by the introduction of the proposed method. These corruptions introduce edgy geometric patterns, and the matching methods find correspondence between them. See Sec. B.1 in the supplementary material for more detailed comparisons.
Robustness against environmental changes.
As the objects to be matched are non-identical, they can appear in largely different environments. The right side of Tab. 1 shows the AUC when only one of the input images is corrupted and the other is clean. As in the left side of Tab. 1, the proposed method boosts the LightGlue and GlueStick for most cases and the average case. Interestingly, the results are better for almost all weather categories. For brightness, it is the only one of all categories that shows a decrease in accuracy, but the drop is slight. Weather categories such as Snow and Frost add strong geometry to the overall image, causing mismatches as occlusions. Heatmap weighting is considered to structurally complement the areas hidden by occlusion. The proposed method makes existing matchers robust to environmental changes in general, with a slight precision reduction in clean images. See Section B.2 in the supplementary material for more detailed comparisons.
5.3 Limitations
We here present failure cases. Figure 7 shows typical cases caused by (a) heatmaps not covering a whole part of objects and (b) great structural differences. In the former case, the importance scores take large only on particular parts of the objects. Further, the large importance scores on the irrelevant object disrupt the matching. This may be resolved by using a better visual explanation model [73, 3, 1, 31, 54]. For the latter case, objects with great structural differences cannot be matched as there is little correspondence between the objects, and eventually, regions of similar colors or textures are matched. To address these cases, one may introduce the vision and language model, such as the segment anything model [24, 44] to complement the visual information with text and part-level mask information.
6 Conclusion
In this study, we addressed a novel task of non-identical object matching, where objects to be matched are only assumed semantically similar, not necessarily identical, as in prior studies. Because of the non-identicality, two objects may be captured in different environments and domains. Thus, this task also encompasses robust matching under image corruptions and domain shifts. Our method exploits object detection and visual explanation as high-level features, thereby strongly focusing existing feature matchers on objects. Our experiments demonstrated the effectiveness of the proposed method in non-identical object matching and relative pose estimation under various image corruptions.
Acknowledgement.
Kazuhiko Kawamoto and Hiroshi Kera were supported by JSPS KAKENHI Grant Numbers JP22H03658 and JP22K17962, respectively. They were also supported by Ricoh Company, Ltd.
Supplementary Material
Appendix A Additional matching results
In Sec. 5.1 of the main text, we evaluated non-identical object matching by inserting our method between SuperPoint [10] and LightGlue [27]. Here, “Ours” in the results refers to the case where only our weighting module is used, without applying the class similarity adapter. The matching results for various image pair show that our method enhances matchers and provides dense and consistent correspondences. To demonstrate this generality for various image pairs, we show more results in three cases: (i) same class (i.e., $y_{A}=y_{B},D_{A}=D_{B}$), (ii) class discrepancy (i.e., $y_{A}\neq y_{B},D_{A}=D_{B}$), (iii) domain shift (i.e., $o_{A}\equiv o_{B}$, $D_{A}\neq D_{B}$). Additionally, we present more results for complex scenes containing multiple objects in a single image.
Setup.
We collected various images with objects from the COCO [26] and ImageNet [9] datasets. To prepare drawing or sketch versions of several images, we used the DreamStudio [^3] DreamStudio: https://beta.dreamstudio.ai/generate, which is backended by the Stable Diffusion [47]. For details of the experiment, refer to Sec. 5.1.
A.1 Success cases
Figure A shows the matching results for the same class case. As in the main text, The introduction of the proposed method turns sparse and inconsistent matching into dense and consistent one. The combination of SuperPoint and LightGlue with our method is robust to object color and shape differences and performs dense and consistent matching. Even in the case with class discrepancy, as shown in Fig. B, the proposed method successfully corresponds the same parts (e.g., eyes to eyes, nose to nose) between various animals. Figure C shows the case of domain shift between images, where our method reduces mismatching caused by image corruptions and image style changes (e.g., photos and drawings). Figure D shows cases where multiple objects are present in an image, demonstrating that our class similarity adapter selects semantically similar object pairs and provides condensed matching between the two objects. These results indicate that the proposed method can perform accurate matching in many cases under the challenging condition of non-identical object matching.
A.2 Failure cases
There are some cases of failure in non-identical object matching, as shown in Fig. E. It is difficult to match objects that are of the same class but have significantly different shapes (e.g., the presence or absence of an airplane propeller) or objects that are similar in color or size to other parts (e.g., eyes and nose that are small, round, and black) of the object. Furthermore, image corruptions sometimes hide the texture of the objects, making it difficult to pinpoint the areas to be matched.
Figure A: Matching results in the same class case. Image pairs show objects of the same class in ImageNet [9] or COCO [26] datasets. (a) A combination of SuperPoint (SP) [10] and LightGlue (LG) [27] only finds a small number of correspondence, and many of them are incorrect. (b) The proposed method significantly improves the matching and performs dense and consistent correspondences.
Figure B: Matching results in the class discrepancy case. Image pairs show objects with class discrepancies in ImageNet [9] or COCO [26] datasets. (a) A combination of SuperPoint (SP) [10] and LightGlue (LG) [27] finds many inconsistent and scattered matches, and it is unclear which is the correct correspondence. (b) The proposed method finds almost all matches for the corresponding parts of the animal, which largely improves the matching to consistency.
Figure C: Matching results in the domain shift case. Image pairs for matching are ImageNet [9] or COCO [26] datasets with common corruptions [20], and illustrations or sketches of a motorcycle, a violin, and a cat are generated by the Stable Diffusion [47]. (a) A combination of SuperPoint (SP) [10] and LightGlue (LG) [27] finds correct matching, but also finds mismatches with many non-correspondence areas. (b) The proposed method reduces these mismatches and further improves to denser matching between objects.
Figure D: Matching results in complex scenes containing multiple objects. (a) A combination of SuperPoint (SP) [10] and LightGlue (LG) [27] causes correspondences to scatter across multiple objects. (b) The proposed class similarity adapter effectively selects semantically similar object pairs and provides condensed and consistent matching between the relevant parts of the objects.
Figure E: Failure cases of non-identical object matching. (a) A combination of SuperPoint (SP) [10] and LightGlue (LG) [27] fails in matching. (b) The proposed method improves several matching but still leaves many corresponding areas unmatched.
Appendix B Robust image matching details
With the same setup in Sec. 5.2 of the main text, we evaluate the effectiveness of the proposed method for LightGlue [27] and GlueStick [38] on the robustness to image corruptions and domain shift in image matching. Here, we consider the maximum angular error at $10\tcdegree$ and $5\tcdegree$ instead of $20\tcdegree$. The results of relative pose estimation with common corruptions [20] on one or both of the input images in the MegaDepth-1500 [56] show that the proposed method improves the robustness of LightGlue and GlueStick. In all experiments, we used the pretrained sparse matchers LightGlue [27], and GlueStick [38], and the dense matchers LoFTR [56] and EfficientLoFTR [64], downloaded from their official repositories. [^4] LightGlue: https://github.com/cvg/LightGlue
GlueStick: https://github.com/cvg/GlueStick
LoFTR: https://github.com/zju3dv/LoFTR
ELoFTR: https://github.com/zju3dv/efficientloftr
B.1 Robustness against common corruptions
In this experiment, common corruptions are added to both input images to demonstrate robustness against corruption. Table A shows the results of relative pose estimation for maximum angular error of $10\tcdegree$ and $5\tcdegree$. As the results show, the proposed method makes the LightGlue and GlueStick robust against most types of common corruptions (roughly, 5% to 10% increase) with a slight decrease in clean accuracy. For the Noise category, LightGlue and GlueStick with our method even surpass dense matchers in both cases.
B.2 Robustness against environmental changes
In contrast to Sec. B.1, in this experiment, common corruptions are added to one of the image pairs to demonstrate robustness against domain shift between images. Table B shows the results of relative pose estimation for maximum angular error of $10\tcdegree$ and $5\tcdegree$. As the results show, the proposed method is robust to LightGlue and GlueStick, and in particular, significantly improves the AUC of the former for all Noise, Blur, and Digital categories, and the latter for all Noise and Blur categories. For the Noise category, LightGlue and GlueStick with our method even exceeds dense matchers in both cases.
Table A: The pose accuracy (AUC) at the maximum angular error of $10\tcdegree$ and $5\tcdegree$ of the relative pose estimation from image pairs MegaDepth-1500 [56] under common corruptions. Both images are corrupted with the same type of corruption. Our method largely improves LightGlue (LG) [27] and GlueStick (GS) [38] for most categories and the average AUC. The results of a dense matchers, LoFTR [56] and Efficient LoFTR (ELoFTR) [64], are shown for reference. Our method improves the robustness of sparse matchers with a marginal decrease in clean accuracy.